You need to make 5 liters of a 12% acetic acid solution. Your lab stocks 5% and 20% acetic acid solutions. How many liters of each stock solution should you mix to make 5 liters of a 12% acetic acid solution?
Answer by mananth(16949)

----------------- percent ---------------- quantity
Acetic Acid type I 5 ----------------x liters
Acetic acid type II 20 ----------------5-x liters
Mixture 12.00% ---------------- 5

5x+20(5-x)=12*5

5x+100- 20x=60
5x-20x= 60-100
-15x=-40
/-15
x=2.67 liters 5.00% Acetic Acid type I
2.33 liters 20.00% Acetic acid type II
m.ananth@hotmail.ca